# DehazeGS: Seeing Through Fog with 3D Gaussian Splatting

### v1

@@ 3.3 DehazeGS
The gradient of depth $d_{\mathcal{G}}(\boldsymbol{x})$ of the Gaussian ellipsoid is detached to prevent gradients from flowing through it. This approach is designed to prevent the gradient from affecting the latent representation of clear Gaussians, while allowing for better optimization of the scattering coefficient. For atmospheric light estimation, it is generally assumed that the atmospheric light in foggy scenes is global. The learnable global atmospheric light parameter $A_{\mathcal{G}}$ applies to all Gaussians. Consequently, the Gaussian distribution in foggy scenes is represented as follows,
@@
$\mathcal{\hat{T}}(\boldsymbol{P})={\sum\limits_{i\in N}t_{\mathcal{G}_{i}}(\boldsymbol{x})}\alpha^{\prime}_{i}{\prod\limits_{j=1}^{i-1}(1-\alpha^{\prime}_{j})},$ (9)
@@
To guide the optimization of Gaussian ellipsoid depths, we utilize DepthAnything V2 Yang et al. (2024b) to predict the depth of foggy images, generating a pseudo-depth map to guide the optimization of the 3DGS-rendered depth. However, when the fog density is high, it can affect the accuracy of depth estimation. To mitigate this issue, we propose incorporating the Contrast Limited Adaptive Histogram Equalization (CLAHE) method. We do not recommend directly using its dehazed image results as supervision, as this could introduce some inherent limitations of the method. Instead, we input their results into a depth estimation network and use the resulting depth map for supervision.
@@ 4.2 Quantitative and Qualitative Results
For the qualitative comparison results, as shown in Fig. 3 and Fig. 4 (we encourage readers to zoom in to further explore the details further), the dehazing results of SeaSplat exhibit noticeable color anomalies, and are missing regions at the edges. Compared to DehazeNeRF, our reconstruction results outperform in both dehazing quality and the details of near and far scenes. Meanwhile, ScatterNeRF fails to reconstruct the scenes on both datasets. Additionally, our model is capable of achieving dehazing reconstruction in both indoor and outdoor complex scenarios. It is worth noting that ”Ours*” in Fig. 3 refers to the results obtained by using pseudo-clear images, generated through novel view synthesis with the original 3DGS, to supervise the potential clear gaussians $\mathcal{G}_{c}$, which is obtained by decoupling from the blurred Gaussian $\mathcal{G}_{f}$. Our model effectively decouples scattering and clear content. When the clear content is further supervised, the dehazing performance of the model can be further improved.

### v2

@@ 3.3 DehazeGS
The depth of the Gaussian ellipsoid $d_{\mathcal{G}}(\boldsymbol{x})$ is normalized to the range $[0,1]$, and its gradient is detached to prevent the gradients from flowing through it. This approach is designed to prevent the gradient from affecting the latent representation of clear Gaussians, while allowing for better optimization of the network. For atmospheric light estimation, it is generally assumed that the atmospheric light in foggy scenes is global. The learnable global atmospheric light parameter $A_{\mathcal{G}}$ applies to all Gaussians. Consequently, the Gaussian distribution in foggy scenes is represented as follows,
@@
$\mathcal{\hat{T}}(\boldsymbol{P})={\sum\limits_{i\in N}t_{\mathcal{G}}(\boldsymbol{x})}\alpha^{\prime}_{i}{\prod\limits_{j=1}^{i-1}(1-\alpha^{\prime}_{j})},$ (9)
@@
To guide the optimization of Gaussian ellipsoid depths, we utilize DepthAnything V2 Yang et al. (2024b) to predict the depth of foggy images, generating a pseudo-depth map to guide the optimization of the 3DGS-rendered depth. We found that, in a very small number of scenarios, introducing the Contrast Limited Adaptive Histogram Equalization (CLAHE) can slightly improve performance. However, in most cases, CLAHE is not required, and directly estimating the depth from multi-view foggy images as supervision is sufficient.
@@ 4.2 Quantitative and Qualitative Results
For the qualitative comparison results, as shown in Fig. 3 and Fig. 4 (we encourage readers to zoom in to further explore the details further), the dehazing results of SeaSplat exhibit noticeable color anomalies, and are missing regions at the edges. Compared to DehazeNeRF, our reconstruction results outperform in both dehazing quality and the details of near and far scenes. Meanwhile, ScatterNeRF fails to reconstruct the scenes on both datasets. Additionally, our model is capable of achieving dehazing reconstruction in both indoor and outdoor complex scenarios.